1. Calculate the present value of $7,000 to be received in 6 years from today if the annual discount rate is 5%.

B) Calculate the future value of $5,500 in 5 years from today if the annual interest rate is 6%.
Answer #1

1.Present value of inflows=cash inflow*Present value of discounting factor(rate%,time period)

=7000/1.05^6

=$5223.51(Approx).

2.We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

A=5500*(1.06)^5

=$7360.24(Approx).